1) Your company currently owns a used pump that has been damaged. A contractor has said that the pump could be repaired at a cost of $35,000, which would make the pump workable for the next five years. If you instead sell the damaged pump for scrap, you can get $5,000 for it. A brand-new pump would cost $80,000, and would last for seven years. If your company's minimum acceptable rate of return is 10% per year before taxes, which option would you recommend?
Answer #1

EUAC of repair = 35000 * (A/P, 10%,5)

= 35000 * 0.263797

= 9232.91

EUAC of new pump = (80000 - 5000)* (A/P, 10%,7)

= 75000 * 0.205405

= 15405.41

As EUAC of repair is less, old pump should be repaired